OpenStudy (anonymous):

Which values of x and y will satisfy y = 5x and 12x + 9y = 60?

OpenStudy (doggy):

y = 5x 12x + 9y = 60 Use substitution and solve the system 12x + 9(5x) = 60 12x + 45x = 60 57x = 60 x=20/19 y = 5(20/19) y= 100/19 i got the answer from here http://openstudy.com/study#/updates/5255709de4b037f05886f028

OpenStudy (doc.brown):

check\[y=5x\]\[\frac{100}{19}=5(\frac{20}{19})\]\[\frac{100}{19}=\frac{100}{19}\]looks good @doggy

OpenStudy (anonymous):

thanks for the help!!!!